OpenStudy (anonymous):

6x - 24 = (5 + 2x + 7)1/2

OpenStudy (anonymous):

@SolomonZelman can you help me ill send you five dollars through the mail if you can help me anwser these questions.

OpenStudy (solomonzelman):

\[\huge\color{red}{6x-24=(5+2x+7)\frac{1}{2}}\]\[\huge\color{red}{6x-24=x+6}\]\[\huge\color{red}{6x-x=6+24}\]\[\huge\color{red}{5x=30}\]\[\huge\color{red}{x=6}\]

OpenStudy (solomonzelman):

You don;t have to send me any money through email, seriously, I can help you for free, just like many people here, this is what the website is for.

OpenStudy (anonymous):

okay thanks your really kind

OpenStudy (solomonzelman):

Anytime, guys!

OpenStudy (anonymous):

Can you do this? @SolomonZelman And yeah you don't send anyone money! This website is for helping. :D ( and for free)

OpenStudy (anonymous):

i know its alot of questions but its cause my math teacher left us a sub and now she wants us to take this final and i dont understand any of it.

OpenStudy (solomonzelman):

Ok, but please post one question per thread, i can try to help you.

OpenStudy (anonymous):

okay sorry its been a while since ive used this website sorry.

OpenStudy (anonymous):

Thank you! @SolomonZelman I really need help with this!

OpenStudy (solomonzelman):

I am really sorry, but can you post the question here, in a thread, so that I won;t have to look here and back into the attachment please?

OpenStudy (anonymous):

What is the length of the midsegment of the trapezoid made by the vertices A(0,5)B(3,3)C(5,-2) and D(-1,2). Show equations and all work that lead up to your answer.

OpenStudy (anonymous):

um ok its alot so ill only as you two so your not overwhelmed@solomanzelman

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (rina.r):

just an advice @Hello_Im_Nick for these types of question which u asked just follow BIMDAS rule then it will be easy for u.

OpenStudy (anonymous):

What's BIMDAS?

OpenStudy (rina.r):

Bracket, Index, Multiplication, Division, Addition and subtraction.

OpenStudy (rina.r):

so u solve ur equation in order of this.

OpenStudy (anonymous):

But can you help me solve this one? please show your steps. I want to do the next one by myself.

OpenStudy (anonymous):

@SolomonZelman i have this final and for the passed week we had a sub and i dont understand anything if someone is willing to help me even though its alot of questions i am willing to mail anyone five dollars for helping me i would offer more but thats all i have? my first question is 2x+3y+3z=5 4x+5y+3z=13 using matrices A.(5,13,13) B. (2,4,3) C. (1,2,-1) D. (1,3,-2) my second question is given the hyperbola (x2/4)-(y^2/9)=1,find asymptotes. A. x=(1/3)y,x=(-1/13)y B.y=(1/3)x,y=(-1/3)x C.y=(3/2)x,y+(-3/2)x D.x=(3/2)y,x=(-3/2)y

OpenStudy (anonymous):

my third what is the equation of the following graph A.x^2+y^2=1 B.x^2-y^2=1 C.x^2+y^2=2 D.x^2+y^2=-2 my fourth question solve the system by graphing y=-x+4(x-3)^2+y^2=25 A.(0,4) and (-3,7) B.(0,5) and (-3,7) C.(4,0) and (-3,7) D.(4,0) (7,-3) my fifth question use the factor theorem to determine (x-1) is a factor of x^3+x^2-16x-16? A. factor b. not factor my sixth solve using matrices 7x+5y=14 4x+3y=9 A.(7,-3) B.(-3,7) C. (3,-7) D.(7,3) My seventh question what is the 3rd term of (a-b)? A.21a^3b^4 B.21a^4b^3 C. 21a^3b^2 D.21a^6b My eighth question is what does the graph x^2-y^2=25 represent? A.circle B.ellipse C.hyperbola D. parabola my nineth question what is the system of equation describes the graph? choices are A.x+3y>or=4 B.x+3y<4 C.x+3y<4 3x-y<2 D.x+3y=4 3x-y>or=2 Reply Using Drawing my tenth question which of the following is the graph of the polynomial f(x)=(x-2)^3? Reply Using Drawing my eleventh question solve using matrices 7x+5y=14 4x+3y=9 A)(7,3) B)(-3,7) C)(3,-7) D)(7,3) My twelveth question is which of the following not a geometric sequence? A)1,2,4,8 B)3,1,-1,-3 C)1/2,1/4,1/8,1/16 D)216,72,24,8

OpenStudy (anonymous):

you dont have to help with all

OpenStudy (solomonzelman):

do you mean use of substitution method to solve the system of linear equations, I can teach you if you want to.

OpenStudy (anonymous):

it asks for elimination for the first question idk how to do that

OpenStudy (rina.r):

really u dont know elimination method do u want me to show u an example.

OpenStudy (anonymous):

sure hopefully ill understand

OpenStudy (anonymous):

Can you guys help me as well....

OpenStudy (rina.r):

Solve the system of equations by elimination 3x-y=5 x-y=3 In this example we will "cancel out" the y term. To do so, we can add the equations together. 3x-y=5 x-y=3 (add equation) 4x=8 Now we can find: x = 2 In order to solve for y, take the value for x and substitute it back into either one of the original equations. x+y=3 2+y=3 y=1 The solution is (x, y) = (2, 1).

OpenStudy (anonymous):

But whats the length of the midsegment?

OpenStudy (rina.r):

go to this link and go through it will help u @Karina1

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

oh im sorry it actually asks for matrices

OpenStudy (rina.r):

what ur question @Hello_Im_Nick ?